A body weighs 0.5kg in air, 0.3kg in water and 0.2kg in a liquid, what is the relative density of the liquid?


1
Expert's answer
2021-01-21T09:56:30-0500

Solution:


The relative (to the water's density) density of the liquid is:


"\dfrac{\rho_l}{\rho_w}=\dfrac{F_l}{F_w}=\dfrac{W_a-W_l}{W_a-W_w}=\dfrac{0.5-0.2}{0.5-0.3}=1.5"


The liquid is 1.5 times denser than water and pushes a body up stronger. Less weight remains for the bottom.


Answer:


The relative density of the liquid is 1.5.